# Double Robustness of Local Projections

### v2

@@ Appendix B Proofs
B.7 Proof of Corollary 4.4
Proposition 4.1 implies that the absolute relative VAR bias $|b_{h}|$ can be made to take any value in $[0,\infty)$ as $\alpha(L)$ varies over the set of all absolutely summable lag polynomials. The corollary then follows from Corollaries 3.1, 3.3 and A.2. ∎
B.8 Proof of Corollary 4.6
The result follows straightforwardly from (4.2) if we can show that the maximal non-centrality parameter equals
$\max_{\alpha(L)\colon\|\alpha(L)\|\leq M}\operatorname{aBias}(\hat{\boldsymbol{\delta}})^{\prime}\operatorname{aVar}(\hat{\boldsymbol{\delta}})^{-1}\operatorname{aBias}(\hat{\boldsymbol{\delta}})=M^{2}\left[\lambda_{\max}(\operatorname{aVar}(\hat{\boldsymbol{\beta}})\operatorname{aVar}(\hat{\boldsymbol{\delta}})^{-1})-1\right].$
But this follows from applying Proposition 4.2 with $R=\operatorname{aVar}(\hat{\boldsymbol{\delta}})^{-1/2}$, since
$\lambda_{\max}\!\left(\operatorname{aVar}(\hat{\boldsymbol{\delta}})^{-1/2}[\operatorname{aVar}(\hat{\boldsymbol{\beta}})-\operatorname{aVar}(\hat{\boldsymbol{\delta}})]\operatorname{aVar}(\hat{\boldsymbol{\delta}})^{-1/2\prime}\right)$
$=\lambda_{\max}\!\left(\operatorname{aVar}(\hat{\boldsymbol{\delta}})^{-1/2}\operatorname{aVar}(\hat{\boldsymbol{\beta}})\operatorname{aVar}(\hat{\boldsymbol{\delta}})^{-1/2\prime}-I_{k}\right)$
$=\lambda_{\max}\!\left(\operatorname{aVar}(\hat{\boldsymbol{\delta}})^{-1/2}\operatorname{aVar}(\hat{\boldsymbol{\beta}})\operatorname{aVar}(\hat{\boldsymbol{\delta}})^{-1/2\prime}\right)-1$
$=\lambda_{\max}\!\left(\operatorname{aVar}(\hat{\boldsymbol{\beta}})\operatorname{aVar}(\hat{\boldsymbol{\delta}})^{-1}\right)-1.\qed$
B.9 Proof of Corollary A.2
We first use Proposition 3.1 to compute $\operatorname{aCov}(\hat{\beta}_{i_{a}^{*},j_{a}^{*},h_{a}},\hat{\beta}_{i_{b}^{*},j_{b}^{*},h_{b}})$. Define $\xi_{j^{*},h,t}=(\xi_{1,j^{*},h,t},\dots,\xi_{n,j^{*},h,t})^{\prime}$ as in Proposition 3.1, but making the dependence on both $i^{*}$ and $j^{*}$ explicit in the notation. Observe that
@@
We now derive $\operatorname{aCov}(\hat{\delta}_{i_{a}^{*},j_{a}^{*},h_{a}},\hat{\delta}_{i_{b}^{*},j_{b}^{*},h_{b}})$ using Proposition 3.2. Observe that the vector process $(\varepsilon_{t}^{\prime}\otimes\tilde{y}_{t-1}^{\prime},\xi_{j_{a}^{*},0,t}^{\prime}\varepsilon_{j_{a}^{*},t},\xi_{j_{b}^{*},0,t}^{\prime}\varepsilon_{j_{b}^{*},t})^{\prime}$ is a martingale difference sequence with respect to the filtration generated by $\{\varepsilon_{t}\}$. Moreover, $E[(\varepsilon_{t}\otimes\tilde{y}_{t-1})\xi_{j^{*},0,t}^{\prime}\varepsilon_{j^{*},t}]=0$ for any $j^{*}$. Hence,
@@
Finally, we compute $\operatorname{aCov}(\hat{\beta}_{i_{a}^{*},j_{a}^{*},h_{a}},\hat{\delta}_{i_{b}^{*},j_{b}^{*},h_{b}})$ using Propositions 3.1 and 3.2. Using arguments similar to above, we obtain

### v3

@@ Appendix B Proofs
B.6 Proof of Corollary˜4.4
Proposition˜4.1 implies that the absolute relative VAR bias $|b_{h}|$ can be made to take any value in $0,\infty)$ as $\alpha(L)$ varies over the set of all absolutely summable lag polynomials. The corollary then follows from Corollaries˜3.1, 3.3 and A.2. ∎
B.7 Proof of Corollary˜A.2
We first use Proposition˜3.1 to compute $\operatorname{aCov}(\hat{\beta}_{i_{a}^{*},j_{a}^{*},h_{a}},\hat{\beta}_{i_{b}^{*},j_{b}^{*},h_{b}})$. Define $\xi_{j^{*},h,t}=(\xi_{1,j^{*},h,t},\dots,\xi_{n,j^{*},h,t})^{\prime}$ as in Proposition˜3.1, but making the dependence on both $i^{*}$ and $j^{*}$ explicit in the notation. Observe that
@@
We now derive $\operatorname{aCov}(\hat{\delta}_{i_{a}^{*},j_{a}^{*},h_{a}},\hat{\delta}_{i_{b}^{*},j_{b}^{*},h_{b}})$ using Proposition˜3.2. Observe that the vector process $(\varepsilon_{t}^{\prime}\otimes\tilde{y}_{t-1}^{\prime},\xi_{j_{a}^{*},0,t}^{\prime}\varepsilon_{j_{a}^{*},t},\xi_{j_{b}^{*},0,t}^{\prime}\varepsilon_{j_{b}^{*},t})^{\prime}$ is a martingale difference sequence with respect to the filtration generated by $\{\varepsilon_{t}\}$. Moreover, $E[(\varepsilon_{t}\otimes\tilde{y}_{t-1})\xi_{j^{*},0,t}^{\prime}\varepsilon_{j^{*},t}]=0$ for any $j^{*}$. Hence,
@@
Finally, we compute $\operatorname{aCov}(\hat{\beta}_{i_{a}^{*},j_{a}^{*},h_{a}},\hat{\delta}_{i_{b}^{*},j_{b}^{*},h_{b}})$ using Propositions˜3.1 and 3.2. Using arguments similar to above, we obtain